OpenStudy (anonymous):

Find the LCM of x^2-10x,x^2+10x I got x^2(x^2+10) Is that correct

OpenStudy (helder_edwin):

let's see \[ \large x^2-10x=x(x-10) \] and \[ \large x^2+10x=x(x+10) \]

OpenStudy (helder_edwin):

so the LCM would be \[ \large x(x+10)(x-10) \]

OpenStudy (anonymous):

Thank you for helping me with this.

OpenStudy (helder_edwin):

u r welcome